OpenStudy (anonymous):

It took Dr.Gru 30 seconds to pump the first $9.00 worth of gasoline into his car,and it took a total of 105 seconds to pump 15 gallons of gas.Assuming gas is pumped at a constant rate, what is the cost for one gallon of gas at this station?

OpenStudy (whpalmer4):

let P be the cost of 1 gallon of gas The pump pumps 15 gallons in 105 seconds = 15 gallons / 105 seconds = 1/7 gal/s In the first 30 seconds, Dr. Gru pumps $9.00 worth of gasoline. In 30 seconds, the pump delivers 1/7 gal/s * 30 s = 30/7 gallons. $9.00 / (30/7 gal) = price per gallon